OpenStudy (anonymous):

Please help Please Part 1: Solve the inequality 2(x + 6) greater than or equal to 3x - 2 and show all of your work. You can use >= to represent the greater than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1. Be sure to include the type of circle and which direction the shading will be.

jigglypuff314 (jigglypuff314):

solve for x... 2(x + 6) >= 3x - 2 so first distribute 2x + 2*6 >= 3x - 2 2x + 12 >= 3x - 2 subtract both sides by 12 -12 -12 2x >= 3x - 14 subtract both sides by 3x -3x -3x -x >= -14 multiply both sides by -1 but remember that when multiplying/dividing by a negative the sign flips so x <= 14 an inequality with an "or equal to" part has a closed circle while one without the "or equal to" part does not